Minder Industries stock has a beta of 1.30. The company just paid a dividend of $.30, and the dividends are expected to grow at 4 percent. The expected return on the market is 13 percent, and Treasury bills are yielding 5.1 percent. The most recent stock price for the company is $68. a. Calculate the cost of equity using the DCF method. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) b. Calculate the cost of equity using the SML method. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

1)

Year 1 dividend = 0.3 (1 + 4%) = 0.312

cost of equity using the DCF method = (D1 / price) + growth rate

cost of equity using the DCF method = (0.312 / 68) + 0.04

cost of equity using the DCF method = 0.004588 + 0.04

cost of equity using the DCF method = 0.0446 or 4.46%

2)

Cost of equity using SML = Risk free rate + beta(market return - risk free rate)

Cost of equity using SML = 0.051 + 1.3(0.13 - 0.051)

Cost of equity using SML = 0.051 + 0.1027

Cost of equity using SML = 0.1537 or 15.37%